# Left-left-right-right magnetic order in spin-1/2 Kitaev-Heisenberg chain

### v1

@@ III Perturbative Luttinger liquid analysis of the Kitaev-Heisenberg model
In this section, we provide an analytical understanding for the origin of the LLRR phase in the spin-1/2 Kitaev-Heisenberg chain based on a perturbative Luttinger liquid analysis. We will work in the $U_{4}$ frame unless otherwise stated. Since the region between $-K$ and $-J$ is dual to the region between FM and $-K$, the perturbative Luttinger liquid analysis can also be applied to explain the FM phase for $\phi\in[\phi_{c2},\pi]$.
@@ III.1 Failure of the classical analysis
in which $\delta=K+2J$. In the region between FM and $-K$ in Fig. 1 (a), we have $J>0$ and $\delta<0$. Therefore, an FM configuration within the $xy$-plane is able to lower the energy in Eq. (20). However, it is clear that the free energy $f$ in Eq. (20) has a U(1) symmetry, corresponding to a rotational symmetry around the $z$-axis. Hence, the classical ground states are infinitely degenerate, since the FM spin direction can point to any direction within the $xy$-plane. This is clearly absurd, since the system only has discrete symmetries and cannot develop a symmetry breaking pattern which has a continuously variable spin orientation.
Quantum fluctuations will break the infinite classical ground state degeneracy. There are two possibilities though. The system either becomes a gapless Luttinger liquid (like the easy-plane XXZ model where the classical analysis also gives an infinite ground state degeneracy but the system is gapless at the quantum level), or is magnetically ordered but with a discrete spontaneous symmetry breaking pattern. We will show that at quantum level, the Luttinger liquid behavior is unstable under the symmetry allowed perturbations, and as a result, the system develops a magnetic order with a discrete symmetry breaking pattern.
@@ III.2 Perturbative Luttinger liquid analysis
The advantage of $V_{2}$ is that the correlation functions of $H^{\prime\prime}_{XXZ}$ in the $xy$-plane are rendered to be dominated by AFM rather than FM fluctuations, which is the commonly used convention in Luttinger liquid theory. After applying the $V_{2}$ transformation, the symmetry operations in Eq. (8) become
@@
in which $S^{\prime\prime+}=S^{\prime\prime x}+iS^{\prime\prime y}$, and $x=na$ ($n\in\mathbb{Z}$) is the spatial coordinate in the continuum limit. The easiest way to treat the perturbation $H_{\delta}$ is to do a first order perturbation, namely, rewriting $H_{LL}$ in terms of the $\theta,\varphi$ fields by replacing the spin operators in $H_{\delta}$ with the expressions in Eq. (30). The first order projection can be evaluated as
@@
The above analysis shows that we need to go beyond first order projection. Since such calculations are difficult, we will not explicitly derive the high order perturbations. Instead, we use a symmetry analysis to figure out the most relevant symmetry allowed terms in the RG sense. It is worth to note that although the symmetry analysis is able to determine the form of the corresponding relevant operator, the signs and order of magnitudes of the coupling constants cannot be figured out. For our purpose, the order of magnitude of the coupling constant is not important, whereas the sign is crucial which determines the specific type of magnetic instability. We will infer the sign of the coupling constant indirectly by numerical calculations in later sections.
@@
in which $g$ is the coupling constant. The operator in Eq. (38) has a scaling dimension $4/\kappa$, which is highly relevant when $\kappa$ becomes large. In fact, $\kappa$ can be arbitrarily large in the vicinity of the FM point in Fig. (1).
@@ III.3 Magnetic ordering
$(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{4h}\rightarrow(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{2}.$ (46)
@@
in which $N_{1}$ is the magnitude of the spin ordering. Rotating back to the original frame by performing $(U_{4})^{-1}$, we obtain
@@
This is a spiral phase in which the spins rotate in the $xy$-plane as the position is moved along the chain.
@@
$\theta^{(\text{II})}_{n}=\frac{n}{2}\sqrt{\pi},\leavevmode\nobreak\ 0\leq n\leq 3,$ (49)
We take one of the four degenerate solutions at $n=1$ as an example, giving $\theta_{0}^{(\text{II})}=\pi/2$. Plugging $\theta_{0}^{(\text{II})}$ into the bosonization formulas in Eq. (30), the spin orientations can be determined as
@@
where $D^{(\text{II})}_{2}\cong D_{2}$.
@@
in which $N_{2}$ is the magnitude of the spin ordering. Rotating back to the original frame by performing $(U_{4})^{-1}$, we obtain
@@
This is a left-left-right-right magnetic order, in which the spins are along the $y$-direction and reverse their directions every two sites.
@@
From the discussions in this section, we see that the sign of the coupling constant $g$ is crucial to determine the magnetic pattern of the system. However, a pure symmetry analysis is not able to give the sign of $g$. The actual magnetic order developed in the system have to be determined by numerics as will be discussed in Sec. VI.
IV LLRR phase in the Kitaev-Heisenberg-Gamma chain

### v2

@@ III Perturbative Luttinger liquid analysis of the Kitaev-Heisenberg model
In this section, we provide an analytical understanding for the origin of the LLRR phase in the spin-1/2 Kitaev-Heisenberg chain based on a perturbative Luttinger liquid analysis. We will work in the $U_{4}$ frame unless otherwise stated. Since the region between $-K$ and $-J$ is dual to the region between FM and $-K$ via the $U_{4}$ transformation, the perturbative Luttinger liquid analysis can also be applied to explain the FM phase for $\phi\in[\phi_{c2},\pi]$.
@@ III.1 Failure of the classical analysis
in which $\delta=K+2J$. In the region between FM and $-K$ on the horizontal axis in Fig. 1, we have $J>0$ and $\delta<0$. Therefore, an FM configuration within the $xy$-plane is able to lower the energy in Eq. (20). However, it is clear that the free energy $f$ in Eq. (20) has a U(1) symmetry, corresponding to a rotational symmetry around the $z$-axis. Hence, the classical ground states are infinitely degenerate, since the FM spin direction can point to any direction within the $xy$-plane. This is clearly absurd, since the system only has discrete symmetries and cannot develop a symmetry breaking pattern which has a continuously variable spin orientation.
Quantum fluctuations will break the infinite classical ground state degeneracy. There are two possibilities though: The system either becomes a gapless Luttinger liquid (like the easy-plane XXZ model where the classical analysis also gives an infinite ground state degeneracy but the system is gapless at the quantum level), or is magnetically ordered but having a discrete spontaneous symmetry breaking pattern. We will show that at quantum level, the Luttinger liquid behavior is unstable under the symmetry allowed perturbations, and as a result, the system develops a magnetic order with a discrete symmetry breaking pattern.
@@ III.2 Perturbative Luttinger liquid analysis
The advantage of applying $V_{2}$ is that the correlation functions of $H^{\prime\prime}_{XXZ}$ in the $xy$-plane are rendered to be dominated by AFM rather than FM fluctuations, which is the commonly used convention in Luttinger liquid theory. After the application of the $V_{2}$ transformation, the symmetry operations in Eq. (8) become
@@
in which $S^{\prime\prime+}=S^{\prime\prime x}+iS^{\prime\prime y}$, and $x=na$ ($n\in\mathbb{Z}$) is the spatial coordinate in the continuum limit.
The easiest way to treat the perturbation $H_{\delta}$ is to do a first order perturbation, namely, rewriting $H_{LL}$ in terms of the $\theta,\varphi$ fields by replacing the spin operators in $H_{\delta}$ with the expressions in Eq. (30). The first order projection can be evaluated as
@@
The above analysis shows that we need to go beyond the first order projection. Since such calculations are difficult, we will not explicitly derive the high order perturbations. Instead, we use a symmetry analysis to figure out the most relevant symmetry allowed terms in the RG sense. It is worth to note that although the symmetry analysis is able to determine the form of the corresponding relevant operator, the signs and order of magnitudes of the coupling constants cannot be figured out. For our purpose, the order of magnitude of the coupling constant is not important, whereas the sign is crucial which determines the specific type of magnetic instability. We will infer the sign of the coupling constant indirectly by numerical calculations in later sections.
@@
in which $g$ is the coupling constant. The operator in Eq. (38) has a scaling dimension $4/\kappa$, which is highly relevant when $\kappa$ becomes large. In fact, $\kappa$ can be arbitrarily large in the vicinity of the FM point on the horizontal axis in Fig. (1).
@@ III.3 Magnetic ordering
$(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{4h}\rightarrow(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{2}^{\text{(I)}}.$ (46)
@@
in which $N_{1}$ is the magnitude of the spin ordering (for derivations, see Appendix B.1). Rotating back to the original frame by performing $(U_{4})^{-1}$, we obtain
@@
This is a spiral phase in which the spins rotate in the $xy$-plane as the position is moved along the chain. The pattern of the magnetic ordering for the other three degenerate ground states can be obtained by applying the broken symmetry transformations to Eq. (48).
@@
$\theta^{(\text{II})}_{n}=\frac{n}{2}\sqrt{\pi},\leavevmode\nobreak\ 0\leq n\leq 3.$ (49)
We take one of the four degenerate solutions at $n=1$ as an example, giving $\theta_{0}^{(\text{II})}=\frac{\pi}{2}$. Plugging $\theta_{0}^{(\text{II})}$ into the bosonization formulas in Eq. (30), the spin orientations can be determined as
@@
where $D^{(\text{II})}_{2}\cong D_{2}$. The symmetry breaking pattern is given by
$(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{4h}\rightarrow(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{2}^{\text{(II)}}.$ (56)
@@
in which $N_{2}$ is the magnitude of the spin ordering (for derivations, see Appendix B.2). Rotating back to the original frame by performing $(U_{4})^{-1}$, we obtain
@@
This is a left-left-right-right magnetic order, in which the spins are along the $y$-direction and reverse their directions every two sites. Similar to the spiral phase, the pattern of the magnetic ordering for the other three degenerate ground states in the LLRR phase can be obtained by applying the broken symmetry transformations to Eq. (58).
@@
From the discussions in this section, we see that the sign of the coupling constant $g$ is crucial to determine the magnetic pattern of the system. However, a pure symmetry analysis is not able to give the sign of $g$. The actual magnetic order developed in the system has to be determined by numerics as will be discussed in Sec. III.4.
III.4 DMRG numerics
We present DMRG simulations to provide numerical evidence for LLRR magnetic order in the Kitaev-Heisenberg chain. For convenience, we stick to the four-sublattice rotated frame in all numerical calculations in this subsection.
III.4.1 Central charge
Figure 5: Fitting of central charge at (a) $\phi=0.38\pi$ and (b) $\phi=0.47\pi$ in the Kitaev-Heisenberg model using the formula in Eq. (59) for open boundary conditions. The central charge is taken as the average value of the fitted slopes. DMRG numerics are performed on an open chain of $L=256$ sites. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1000$, $\epsilon=10^{-11}$.
When the low energy physics of the system is described by a critical theory, the entanglement entropy $S_{L}(x)$ for a subregion of length $x$ in a finite chain of length $L$ is predicted by conformal field theory (CFT) to scale as Calabrese2009 ; Ejima2011
$S_{L}(x)=\lambda\frac{c}{3}\ln\left[\frac{L}{\pi}\sin\left(\frac{\pi x}{L}\right)\right]+\cdots,$ (59)
in which $c$ is the central charge, $\lambda=1$ (and $1/2$) for periodic (and open) boundary conditions Laflorencie2006 , and $\cdots$” represents subleading terms. According to CFT, Luttinger liquids have central charge $c=1$, whereas the smallest possible central charge is $c=0.5$ which is the value for Ising critical theory DiFrancesco1997 . Hence, whether a reliable central charge $c$ can be extracted and its fitted value can be used as criteria to tell if the system is critical or not.
Fig. 5 (a) and (b) show the fitting of the central charge values at two representative points $\phi=0.38\pi$ and $\phi=0.47\pi$, respectively, using the formula in Eq. (59) for open boundary conditions. The entanglement entropy in both Fig. 5 (a) and (b) exhibit an oscillating behavior, and the values of the central charge are taken as the average of the fitted slopes of the red and blue lines in each figure. It is clear from Fig. 5 (a,b) that while the data for $\phi=0.48\pi$ is consistent with a $c=1$ behavior, the low energy physics of the point of $\phi=0.37\pi$ is not described by Luttinger liquid theory.
To investigate the behavior of the central charge in the whole parameter region, Fig. 2 shows the fitted values of the central charge in the region $\phi\in(\arctan(2),\pi/2)$ (where $J=\cos(\phi)$, $K=-\sin(\phi)$) extracted from DMRG numerical results for entanglement entropy at different system sizes with open boundary conditions, using the formula in Eq. (59) and the same fitting method as the one used for Fig. 5. It can observed from Fig. 2 that when $\phi<0.4\pi$, the central charge value decreases by increasing system size, having a tendency to approach zero in the limit of infinite sizes. This is consistent with the analytic prediction in Sec. III on the existence of a magnetically ordered phase when $\phi$ is close to $\arctan(2)$. On the other hand, when $\phi>0.4\pi$, the value of the numerically extracted central charge increases by increasing the system size, tending to approach the value of $1$ in the large size limit, thereby indicating a $c=1$ gapless phase in the region $\phi\in(0.4\pi,0.5\pi)$. The low energy field theory describing this gapless phase is non-perturbative in nature, which remains unclear and will be left for future studies. Also notice that $\phi=0.4\pi$ is special in Fig. 2 in the sense that the value of $c$ remains nearly the same for all the three system sizes. Whether the low energy physics at $\phi=0.4\pi$ is described by one of the minimal models in CFT DiFrancesco1997 remains an open question, and will be left for future more detailed investigations.
Figure 6: Spin correlation functions $\langle S^{\alpha}_{1}S^{\alpha}_{x}\rangle$ in the $U_{4}$ frame versus $\sin(\pi r/L)$ on a log-log scale in the Kitaev-Heisenberg model for (a) $\alpha=x,y$ at $\phi=0.37\pi$, (b) $\alpha=z$ at $\phi=0.37\pi$, (c) $\alpha=x,y$ at $\phi=0.48\pi$, and (d) $\alpha=z$ at $\phi=0.48\pi$. DMRG numerics are performed on systems of $L=96$ sites using periodic boundary conditions. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1200$, $\epsilon=10^{-10}$.
III.4.2 Correlation functions
To further distinguish the $\phi<0.4\pi$ region from the $\phi>0.4\pi$ region, we calculate and compare spin correlation functions $\langle S_{1}^{\prime\alpha}S_{x}^{\prime\alpha}\rangle$ in the $U_{4}$ frame ($\alpha=x,y,z$) in the two regions. Notice that if the system is described by the Luttinger liquid theory, then the exponent for transverse correlation functions ($\alpha=x,y$) is predicted to be $1/(2\kappa)$, and the exponent for the longitudinal one ($\alpha=z$) is $2\kappa$ The product of the values of these two types of exponents is equal to $1$.
Fig. 6 (a,b) show the transverse and longitudinal correlation functions at $\phi=0.37\pi$, respectively. As is clear from Fig. 6 (a,b), the exponents of transverse correlation functions at $\phi=0.37\pi$ are nearly equal (approximately $0.147$) and very small, indicating a slowly decaying correlation behavior. In particular, the product of transverse and longitudinal exponents is equal to $0.33$, far less than $1$, indicating a non-Luttinger liquid behavior, consistent with the perturbative Luttinger liquid analysis in Sec. III.2. We note that in the magnetically ordered LLRR phase, the transverse correlation functions should asymptotically approach non-vanishing constant values in the long distance limit, which seems to be not consistent with the decay behavior in Fig. 6 (a,b). However, such discrepancy is most possibly a finite size artifact. Notice that the term $g\cos(4\sqrt{\pi}\theta)$ (see Eq. (38)) leading to the LLRR order is a high order effect, not appearing in the first order perturbation. Hence the effect of $g\cos(4\sqrt{\pi}\theta)$ only becomes prominent after a long “time” of RG flow, meaning that the system size has to be very large. For moderate system sizes available in numerical calculations like the one chosen in Fig. 6, the length scale may be not long enough to observe the correct behavior for magnetic orderings.
Figure 7: Spin expectation values in the Kitaev-Heisenberg model in the $U_{4}$ frame throughout the chain at $\phi=0.36\pi$ in the presence of a small field of strength $2\times 10^{-4}$ along (a) $(1,1,0)$-direction, and (b) $(0,1,0)$-direction. DMRG numerics are performed on a system of $L=96$ sites with periodic boundary conditions. In (a), the data points for spin expectations values along $x$- and $y$-directions coincide, whereas in (b), those along $x$- and $z$-directions coincide. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1200$, $\epsilon=10^{-10}$.
Fig. 6 (c,d) show the correlation functions at $\phi=0.48\pi$. The product of exponents for the transverse and longitudinal correlations is approximately $1.1$, which is very close to $1$. In addition, the central charge value at $\phi=0.48\pi$ obtained in Fig. 2 is close to $1$. Both the exponents for correlation functions and the central charge value are consistent with Luttinger liquid theory. In addition, exponents for transverse correlations are predicted to be $0.5$ for free fermion systems, which is close to the numerically obtained values in Fig. 6 (c). Therefore, existing DMRG numerics at $\phi=0.48\pi$ are consistent with a low energy theory of a free fermion. An analytical understanding for the $\phi>0.4\pi$ region is still lacking and will be left for future considerations.
III.4.3 Magnetic ordering in the LLRR phase
Notice that there are two possible types of magnetic orders according to the analysis in Sec. III and Sec. IV, namely, the spiral order and the LLRR order. We will study the response of the spin expectation values to a small uniform external magnetic field along certain directions, where the direction in the $U_{4}$ frame can be chosen as $\frac{1}{\sqrt{2}}(\hat{x}+\hat{y})$ and $\hat{y}$ for the spiral and LLRR orders, respectively. The added terms $H_{xy}$ and $H_{y}$ in the Hamiltonian for the $\frac{1}{\sqrt{2}}(\hat{x}+\hat{y})$ and $\hat{y}$ fields in the $U_{4}$ frame are given by
$H_{xy}$ $=$ $-\frac{1}{\sqrt{2}}h_{xy}\sum_{i}(S_{i}^{\prime x}+S_{i}^{\prime y}),$
$H_{y}$ $=$ $-h_{y}\sum_{i}S_{i}^{\prime y}.$ (60)
When there is spontaneous symmetry breaking, it is expected that an infinitesimal field is enough to produce finite spin expectation values in the thermodynamic limit.
We perform DMRG numerics in the four-sublattice rotated frame, and accordingly, the patterns of spin expectation values of the spiral phase and the LLRR phase in the $U_{4}$ frame are given by Eq. (47) and Eq. (57), respectively. When a uniform $(1,1,0)$-field (or $(0,1,0)$-field) is applied in the $U_{4}$ frame, it is expected that the spin pattern in Eq. (47) (or Eq. (57)) is developed. Fig. 7 (a,b) displays the numerical results of the spin expectation values at $\phi=0.36\pi$ when a $h=2\times 10^{-4}$ field is applied along the $(1,1,0)$- and $(0,1,0)$-directions, respectively. As can be clearly seen, the pattern in Fig. 7 (a) is consistent with Eq. (47), and the one in Fig. 7 (b) is consistent with Eq. (57).
However, there arises the question which one of the two orders applies to the system. Before answering this question in Sec. III.4.4, we explain why both responses have the correct pattern expected from the corresponding magnetic ordering, or more precisely, why the system has the response consistent with the spiral order when the actual order is LLRR, and vice versa. In particular, notice that because of this subtlety, the method for calculating the spin expectation values in the presence of a thermodynamically infinitesimal field is not able to distinguish between the two orders.
Suppose the actual magnetic order of the system is the LLRR order. As discussed in Eq. (39), the ground states are four-fold degenerate in the thermodynamic limit, which can be related by each other by the broken symmetry transformations. Denote $|G_{1}\rangle$ as the ground state having spin expectation values given in Eq. (47). Then the other three ground states $|G_{i}\rangle$ ($i=2,3,4$) in the four-sublattice rotated frame are given by
$|G_{2}\rangle$ $=$ $R(\hat{z},\frac{\pi}{2})T_{a}|G_{1}\rangle$
$|G_{3}\rangle$ $=$ $[R(\hat{z},\frac{\pi}{2})T_{a}]^{2}|G_{1}\rangle$
$|G_{4}\rangle$ $=$ $[R(\hat{z},\frac{\pi}{2})T_{a}]^{3}|G_{1}\rangle,$ (61)
where the broken symmetry operations are chosen as $[R(\hat{z},\frac{\pi}{2})T_{a}]^{i}$ ($i=2,3,4$). Clearly, the expectation values of spin operators are along $\hat{y}$-, $-\hat{x}$-, $-\hat{y}$-, and $\hat{x}$-directions for $|G_{1}\rangle$, $|G_{2}\rangle$, $|G_{3}\rangle$, and $|G_{4}\rangle$, respectively. As a result, when a small field is applied along the $(1,1,0)$-directions, the energies of the states $|G_{1}\rangle$ and $|G_{4}\rangle$ will be lowered by the same amount, whereas the energies of $|G_{2}\rangle$ and $|G_{3}\rangle$ will be raised. Hence, any linear combination of $|G_{1}\rangle$ and $|G_{4}\rangle$ is energetically favored in the thermodynamic limit, though there can be small energy differences in finite systems.
Now consider the state $|\Psi\rangle$ defined as
$|\Psi\rangle=\frac{1}{\sqrt{2}}(|G_{1}\rangle+|G_{4}\rangle).$ (62)
The state $|\Psi\rangle$ is among the degenerate manifold of states when a small field in the $(1,1,0)$-direction is applied. On the other hand, as can be easily checked, the spin expectation values in the state $|\Psi\rangle$ have exactly the same pattern as Eq. (47), which is the pattern supposed to be the one for the spiral order in the $U_{4}$ frame. This explains why the system has a response to a $(1,1,0)$-field as if it has a sprial order, although the actual order is LLRR.
III.4.4 Energy-field relation in the LLRR phase
To further confirm the existence of magnetic orders in the LLRR phase, we calculate ground state energy as a function of applied small fields at $\phi=0.36\pi$. When the system is magnetically ordered, the energy change is expected to be linear with respect to the applied field in the thermodynamic limit. Fig. 8 shows the DMRG numerical results for ground state energy per site $E_{0}/L$ as a function of the applied small fields along $(1,1,0)$- and $(0,1,0)$-directions for two system sizes $L=48$ and $96$. It is clear that the relation between energy change and applied fields are linear in Fig. 8, confirming the existence of a magnetic order.
Figure 8: Ground state energy per site $E_{0}/L$ in the Kitaev-Heisenberg model as a function of applied small fields $h_{xy}$ (dashed lines) and $h_{y}$ (dashed lines), where $h_{xy}$ and $h_{y}$ are along $(1,1,0)$- and $(0,1,0)$-directions, respectively. DMRG numerics are performed on system sizes $L=48$ (blue lines) and $L=96$ (red lines) at $\phi=0.36\pi$ using periodic boundary conditions. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1200$, $\epsilon=10^{-11}$.
Furthermore, as a byproduct of Fig. 8, it can be observed that under the same field strength, the energy is lowered more by the $h_{y}$ field than the $h_{xy}$ field, which indicates that the actual magnetic order is the LLRR order (given by Eq. (57) in the $U_{4}$ frame), rather than the spiral order (Eq. (47) in the $U_{4}$ frame). In fact, suppose the magnetic order is determined by Eq. (57). Then according to Eq. (61), we have
$\langle G_{1}|\vec{S}_{j}|G_{1}\rangle$ $=$ $N_{2}(0,1,0)^{T},$
$\langle G_{4}|\vec{S}_{j}|G_{4}\rangle$ $=$ $N_{2}(1,0,0)^{T},$ (63)
hence
$\frac{1}{L}\langle G_{1}|H_{y}|G_{1}\rangle$ $=$ $-N_{2}h_{y}$
$\frac{1}{L}\langle\Psi|H_{xy}|\Psi\rangle$ $=$ $-\frac{1}{\sqrt{2}}N_{2}h_{xy},$ (64)
where $H_{y}$ and $H_{xy}$ are defined in Eq. (60), and $|\Psi\rangle$ is defined in Eq. (62). Clearly, if $h_{y}=h_{xy}=h$, then the energy change $\Delta E_{xy}=\frac{1}{L}\langle\Psi|H_{xy}|\Psi\rangle$ by the $h_{xy}$ field compared with the zero field case is smaller by a factor of $\frac{1}{\sqrt{2}}$ than the energy change $\Delta E_{y}=\frac{1}{L}\langle G_{1}|H_{y}|G_{1}\rangle$ by the $h_{y}$ field. As can be inspected in Fig. 8, $\Delta E_{y}$ is indeed larger than $\Delta E_{xy}$ when $h_{y}=h_{xy}$, consistent with an order along $(0,1,0)$-direction. However, the ratio $\Delta E_{xy}/\Delta E_{y}$ deviates from $\frac{1}{\sqrt{2}}$ significantly. This is possibly a finite size effect, since there is still reminiscence of the Luttinger liquid behavior when the system size is not large enough.
III.4.5 Excitation gap in the $c=1$ phase
To further confirm the gapless nature of the $\phi>0.4\pi$ region, we numerically calculate the excitation gap at $\phi=0.48\pi$ to verify the $1/L$ scaling of the gap. As shown in Fig. 9, the excitation gap is clearly linear as a function $1/L$. The extrapolation of the gap to the $L\rightarrow\infty$ limit gives a gap value of $2\times 10^{-4}$, which is approximately zero, consistent with a gapless behavior.
Figure 9: Excitation gap vs. $1/L$ in the Kitaev-Heisenberg model at $\phi=0.48\pi$. DMRG numerics are performed on system sizes $L=24,48,72,96$ under periodic boundary conditions. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1200$, $\epsilon=10^{-10}$.
IV Distorted-LLRR phase in Kitaev-Heisenberg-Gamma chain